Butane is used as a fuel in camping gas stoves.100 cm3 of butane is burned in excess oxygen .calculate the volume of carbon dioxide produced at r.t.p
1
Expert's answer
2017-11-01T14:01:07-0400
2 C4H10 + 13 O2 = 8 CO2 + 10 H2O
n = V/22.4
n (C4H10) = 0.1/22.4 = 0.0047 mol
n (CO2) = 4 · n (C4H10) = 4 · 0.0047 = 0.018 mol
V (CO2) = n · 22.4 = 0.018 · 22.4 = 0.403 L = 403 cm3